A mixture of 90.0 grams of CH4 and 10.0 grams of argon has a pressure of 250 torr under conditions of constant temperature and volume. The partial pressure of CH4 in torr is:


1
Expert's answer
2021-11-05T00:57:08-0400

Molar mass of CH4 = 16.04 g/mol

Molar mass of argon = 39.9g/mol

From the given masses we can calculate number of moles followed by mole fraction of each gas.

Number of moles of CH4 = 90.0/16.04 = 5.61

Number of moles of argon= 10/39.9 = 0.25

Mole fraction of CH4 = 5.61/(5.61+0.25)= 0.95

Hence partial pressure= Mole fraction × Total pressure

= 0.95× 250 torr

= 239 torr